OpenStudy (anonymous):

There are 25 multiple choice questions. 6 points for a correct answer,2 points for not answering a question, and 1 point for a wrong answer. billy got x right answers and y questions unawswered.write an expression in simplified form for billys score

OpenStudy (anonymous):

Answer should be 5x+y+25but i dont know how to get it

OpenStudy (anonymous):

But my book says 5x+y+25

OpenStudy (anonymous):

25-x-y=z 6x+2y+z=score

OpenStudy (anonymous):

6x+2y+(25-x-y)

OpenStudy (anonymous):

5x+y+25=score

OpenStudy (anonymous):

Please show steps

OpenStudy (anonymous):

see how I substituted 25-x-y in for z in the second equation. Then i just simplified

OpenStudy (anonymous):

The question is worded kinda funny. Its wierd to think of him getting points for a wrong answer and leaving it blank. But that is how the book did it.

OpenStudy (anonymous):

Thanks :)